OpenStudy (anonymous):

If x^3+y^3=26, find the value of d^2y/dx^2 at the point (-1,3). The value at d^2y/dx^2 at the point (−1,3) is

OpenStudy (bradely):

x^3+y^3=26 Take derivative both sides 3x^2+3y^2dy/dx = 0 dy/dx =-x^2/y^2 d^2y/dx^2 =(y^2(2x)-2xdy/dx)/y^4 =(2xy^2-2x(-x^2/y^2))/y^4 =(2xy^2+2x^4/y^2)/y^4 Substitute x=-1 and y=3 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

Thanks.